On January 1 Patti made a $90,000 interest-free loan to her son, Alex, who used the money to pay off the $40,000 mortgage on his personal residence and also to buy a $50,000 certificate of deposit. Alexs only income for the year is his salary of $45,000 and $2,000 interest income on the certificate of deposit. The relevant Federal interest rate is 3% compounded semiannually. The loan is outstanding for the entire year. 5. Based on the above information, what is the effect of this loan on Pattis gross income for the year? a.
Answer #1

It is given that Patti has given an interest loan of $ 90000 to his Son due to which he is loosing an Interest amount @ 3% (i.e specified by federal interest semi annualy)

Calculation of interest for the year

A = P(1+r/n)nt

=$ 90,000( 1 + 3/2)2*1

=$ 92,720

Interest =Amount - Principal=92,720 - 90,000 = 2,720

Hence he is loosing an amount of $ 2720 as he is givig interest free loan to his son